If y(1) = – 3, y(3) = 9, y(4) = 30, and y(6) = 132, find the four-point

Lagrange interpolation polynomial that takes the same values as the function y at the given points.


1
Expert's answer
2021-01-04T20:19:29-0500

"\def\arraystretch{1.5}
 \begin{array}{c:c:c:c:c}
 x & 1 & 3 & 4 & 6 \\ \hline
 y & -3 & 9 & 30 & 132
\end{array}"

The Lagrange interpolation polynomial is

"f(x)= \frac{(x-x_1)(x-x_2)(x-x_3)}{(x_0-x_1)(x_0-x_2)(x_0-x_3)}y_0+ \frac{(x-x_0)(x-x_2)(x-x_3)}{(x_1-x_0)(x_1-x_2)(x_1-x_3)} y_1+ \frac{(x-x_0)(x-x_1)(x-x_3)}{(x_2-x_0)(x_2-x_1)(x_2-x_3)}y_2+ \frac{(x-x_0)(x-x_1)(x-x_2)}{(x_3-x_0)(x_3-x_1)(x_3-x_2)} y_3=\\"

"=\frac{1}{10}(x-3)(x-4)(x-6)+\frac{3}{2}(x-1)(x-4)(x-6)-5(x-1)(x-3)(x-6)+\frac{22}{5}(x-1)(x-3)(x-4)=\\"

"\frac{1}{10}(x^3-13x^2+54x-72)+\frac{3}{2}(x^3-11x^2+34x-24)-5(x^3-10x^2+27x-18)+\frac{22}{5}(x^3-8x^2+19x-12)=\\"

"=x^3-3x^2+5x-6" .


Answer: "f(x)=x^3-3x^2+5x-6."